Consider an economy where only three types of vehicles are produced and sold, namely: Tesla, GM and BMW. The following table shows the quantity sold (in millions) and the price per unit of each vehicle: 2011 2012 40 $4 200 2013 50 $4.50 210 $4 13 $7 2014 Tesla GM BMW Quantity produced and sold 40 Price per unit Quantity produced and sold200 Price per unit Quantity produced and sold12 Price per unit $3 $2$3.75 S4 65 $5 300 2015 70 $5 375 $2.50$2.50 17 $7 20 $7 $6 Copy and paste the above table in excel to conduct the computations for each question Present your results clearly in a table as they apply to each questiona) Compute the yearly total value for each vehicle and the yearly nominal GDP for the economy. What is the growth rate in nominal GDP for each year?
Answer #1
Telsa GM BMW
Price($) Quatity Yearly total Value( in million $) = P*Q Price($) Quatity Yearly total Value( in million $) = P*Q Price($) Quatity Yearly total Value( in million $) = P*Q Nominal GDP(in million $)
2011 3 40 120 2 200 400 4 12 48 568
2012 4 40 160 3.75 200 750 6 12 72 982
2013 4.50 50 225 4 210 840 7 13 91 1156
2014 5 65 325 2.50 300 750 7 17 119 1194
2015 5 70 350 2.50 375 937.50 7 20 140 1427.50

NGDP Growth rate from 2011 to 2012 = [(982 - 568) / 568] * 100 = 72.89%

NGDP Growth rate from 2012 to 2013 = [(1155 - 982) / 982] * 100 = 17.72%

NGDP Growth rate from 2013 to 2014 = [(1194 - 1156) / 1156] * 100 = 3.29%

NGDP Growth rate from 2014 to 2015 = [(1427.50 - 1194) / 1194] * 100 = 19.56%